OpenStudy (luigi0210):

Evaluate the integral:

OpenStudy (luigi0210):

\[\LARGE \int sin^2(\pi x)~cos^5(\pi x)~dx\]

OpenStudy (anonymous):

assume cos(πx) =z and proceed further

ganeshie8 (ganeshie8):

\[\int sin^2(\pi x)~cos^5(\pi x)~dx = \int sin^2(\pi x)~\left(1-sin^2(\pi x)\right)^2 \cos (\pi x)~dx\]

ganeshie8 (ganeshie8):

sin(πx) =z will work..

OpenStudy (anonymous):

\[\frac{ 1 }{ \pi }\int\limits \sin^2(u)\cos^5(u)du \] u = pi*x , du = pi dx \[\frac{ 1 }{ \pi } \int\limits \sin^2(u)(1-\sin^2(u))^2\cos(u)du\] sin^2u = 1-cos^2u \[\frac{ 1 }{ \pi } t^2(1-t^2)^2dt\] t = sinu, dt = cosudu etc, etc.

OpenStudy (anonymous):

there should be a integral sign before t^2 lol

OpenStudy (anonymous):

A good integration problem , although this isn't rocket science

OpenStudy (anonymous):

\[\frac{ \sin ^{3}(\pi x) (15\cos ^{4}(\pi x ) + 12\cos ^{2}(\pi x)+8)}{ 105\pi } + C\]

OpenStudy (anonymous):

This must be the answer

OpenStudy (anonymous):

Is this the answer

OpenStudy (anonymous):

Substitute u =pi x dx =[1/pi] du Henceforth it would be easy , You will need to apply integral substitution once more :)

OpenStudy (anonymous):

THANK YOU MICHIO THANK YOU SO MUCH FOR YOUR COMMITMENT

OpenStudy (anonymous):

And Apply integral reduction twice before that

OpenStudy (luigi0210):

Okay, thank you all :D

OpenStudy (anonymous):

Nice integration problem !

OpenStudy (anonymous):

Good luck!

OpenStudy (ikram002p):

>.< michio